OpenStudy (anonymous):

WILL MEDAL, HELP IF YOU'RE INTERESTED? (: The first four terms of a sequence are shown below: 8, 5, 2, -1 Which of the following functions best defines this sequence? A) f(1) = 8, f(n+1) = f(n) + 3; for n greater than or equal to 1 B) f(1) = 8, f(n+1) = f(n) - 5; for n greater than or equal to 1 C) f(1) = 8, f(n+1) = f(n) + 5; for n greater than or equal to 1 D) f(1) = 8, f(n+1) = f(n) -3; for n greater than or equal to 1

OpenStudy (rock_mit182):

d

OpenStudy (rock_mit182):

5-8 = -3

OpenStudy (rock_mit182):

2-5= -3

OpenStudy (rock_mit182):

-1 -2 = -3 again so if you realize the pattern you can conclude that the answer is d

OpenStudy (anonymous):

Thank you so so much! Do you mind helping me with some more?

OpenStudy (rock_mit182):

sure but open another post